OpenStudy (anonymous):

solve each system by substitution y=-4x-12 y=-3x-10

OpenStudy (anonymous):

\[-4x-12=-3x-10\] solve for \(x\)

OpenStudy (anonymous):

add \(4x\) to both sides to get \[-12=x-10\] then add \(10\) to get \[-2=x\]

OpenStudy (anonymous):

if \(x=-2\) then \(y=-4(-2)-12=8-12=-4\) so your answer is \((-2,-4)\)

OpenStudy (anonymous):

thanks that helped alot

OpenStudy (anonymous):

yw